# Unsupervised Model Learning for Quantum Impurity Systems

### v1

@@ IV Quantifying the learning progress
While energy scales such as $T_{K}$ and the free energy are invariant under the Wilsonian RG, thermodynamic observables flow under the RG transformation. To understand how observables flow it is instructive to interpret the RG procedure as a reparametrization of the free energy. In this picture the renormalization is achieved by replacing the bare coordinate system with a new coordinate system that leaves $\mathcal{F}$ invariant $\mathbf{C}\rightarrow\mathbf{C}^{\prime}(\mathbf{C})$ (for simplicity we drop the explicit dependence of the renormalized parameters on the bare parameters from the notation). In this picture thermodynamic observables can be interpreted as covariant tensors that can be transformed as such
Figure 2: Impurity-bath spin-spin expectation value ${\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle}$ at zero temperature for the sAIM Eq. 22 (blue line) compared with the corresponding value in the minimal model Eq. 24, optimized by F-learning (green). The inset shows their ratio, which captures the different renormalizations along the two paths to the same SC fixed point. Plotted as a function of $U_{d}$ for $0.38V^{2}/U_{d}=$ 0.3.
between coordinate systems [61]
$\langle\hat{h}_{i}\rangle_{\rm eff}$ $=\frac{\partial}{\partial\theta_{i}}\mathcal{F}(\boldsymbol{\theta})$ (43)
$=\sum_{j}\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}\frac{\partial}{\partial\theta^{*}_{j}}\mathcal{F}(\boldsymbol{\theta}^{*})=\sum_{j}\chi^{*}_{\theta_{j}}\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}\;,$
with $\boldsymbol{\theta}^{*}$ being the FP couplings and $\chi^{*}_{\theta_{i}}$ the FP expectation value. The matrix $\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}$ can be obtained from the $\beta$-function (Eq. 19) and allows us to make a connection between the FP value of an observable and the unrenormalized expectation values. Note that we have replaced $\mathbf{C}$ with $\boldsymbol{\theta}$ to emphasize the effective model parameters (Eq. 15). Eq. 43 shows that the bare expectation value is a linear combination of all possible FP expectation values, weighted by the dependence of the FP parameter on the bare parameters.
As an example we consider the Kondo model, for which it is well known that the SC FP is reached for a renormalized coupling of $J^{*}=\mathcal{O}(1)$. Therefore, the electrons in the system lock into a singlet to screen the impurity spin. The impurity-bath spin correlator ${\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle}=\chi^{*}_{J}$ athe FP follows as:
$\partial_{J}\mathcal{F}(J^{*})=\chi^{*}_{J}=-3/4\;.$
Despite that the same mechanism dominates the bare Kondo model, one does not measure the FP value $\chi^{*}_{J}$ but
$\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle=\chi^{*}_{J}\frac{\partial J^{*}}{\partial J}\;.$
The scaling laws of the Kondo model are known analytically from several different approaches [64], however only to a certain order in perturbation theory and therefore the estimation for $\partial_{J}J^{*}$ vanishes as $D\rightarrow 0$. This prohibits the exact analytic calculation of ${\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle}$ from $\chi^{*}_{J}$ (the true scaling of bare couplings is far beyond tractability).
Despite that, it is still insightful to express the expectation values of effective interactions $\hat{h}_{i}$ (or $\hat{P}_{i}$) in the bare model in this formulation, even if in the end we evaluate them numerically. Because the effective interactions do not occur in the bare Hamiltonian they have to be added by means of source terms, which we also denote with $\theta_{i}$. The thermodynamic expectation value can now be calculated by differentiating the free energy with respect to the source term
$\langle\hat{h}_{i}\rangle_{\rm bare}$ $=\frac{\partial}{\partial\theta_{i}}\mathcal{F}(\mathbf{C};\theta_{i})\Big{|}_{\theta_{i}=0}$ (44)
$=\left[\frac{\partial\theta^{*}_{i}}{\partial\theta_{i}}\frac{\partial}{\partial\theta^{*}_{i}}+\sum_{j}\frac{\partial C^{*}_{j}}{\partial\theta_{i}}\frac{\partial}{\partial C^{*}_{j}}\right]\mathcal{F}(\mathbf{C}^{*};\theta^{*}_{i})\Big{\rvert}_{\theta_{i}=0}$
where $\mathbf{C}^{*}$ are the bare FP parameters. The effective model is chosen such that as $D\rightarrow 0$ it approaches the same FP as the bare model. Thus, the bare FP parametrization involves the same interactions and parameters as the effective model
$\langle\hat{h}_{i}\rangle_{\rm bare}=\frac{\partial}{\partial\theta_{i}}\mathcal{F}(\mathbf{C};\theta_{i})\Big{|}_{\theta_{i}=0}=\sum_{j}\chi^{*}_{\theta_{j}}\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}\Big{|}_{\theta_{i}=0}\;,$ (45)
however here we have $\theta^{*}_{j}({\mathbf{C}})$, while in we have Eq. 43 $\theta^{*}_{j}({\boldsymbol{\theta}})$. The FP expectation values $\chi^{*}_{\theta_{j}}$ are identical for bare and effective models, but the bare expectation value differs from the effective one only by the different dependence of the FP parameters on the source term. Therefore, matching observables ($\langle\hat{h}_{i}\rangle_{\rm bare}=\langle\hat{h}_{i}\rangle_{\rm eff}$) actually means that
$\Bigg{[}\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}\Bigg{]}_{\rm eff}=\Bigg{[}\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}\Big{|}_{\theta_{i}=0}\Bigg{]}_{\rm bare}\;,$ (46)
is matched between bare and effective model. If both models would flow in the same manner to the FP this would imply that observables and energy scales can be matched at the same time. However, the way in which bare and effective model flow to the FP is not in general the same. The bare RG-flow typically involves more FPs than the effective RG-flow, which by construction is simpler. For instance, the sAIM visits the free orbital FP, which the Kondo model does not. As a consequence when bare and effective observables coincide, the free energies are not necessarily agreeing, and vice versa.
For the case of the sAIM and the Kondo model Eq. 24 acting as minimal model, it is found that the expectation value of the effective interaction in the effective model is upper bounded by the expectation value in the bare model
$\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle_{\rm eff}\leq\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle_{\rm bare}\;.$ (47)
We demonstrate this fact for a wide range of sAIM parameters in Fig. 2. We can rewrite this upper bound using the FP expectation value and the FP coupling derivative as
$\chi^{*}_{J}\left[\frac{\partial J^{*}}{\partial J}\right]_{\rm eff}\leq\chi^{*}_{J}\left[\frac{\partial J^{*}}{\partial J}\right]_{\rm bare}$ (48)
and using $\chi^{*}_{J}=-\tfrac{3}{4}$ if follows immediately that
$\left[\frac{\partial J^{*}}{\partial J}\right]_{\rm bare}\Big{/}\left[\frac{\partial J^{*}}{\partial J}\right]_{\rm eff}\leq 1\;,$ (49)
which is exactly what we observe in Fig. 2 and also for models beyond the minimal model. In the following we show how this discrepancy affects the ability of various effective models to capture the Kondo temperature and how the approximation can be improved by including RG-irrelevant terms.
In equation Eq. 30 we have demonstrated the importance or relating the identity operator of the effective impurity Fock space to the bare impurity Fock space. However, the identity operator is special in the sense that it does not obey Eq. 43
$\langle\hat{\Omega}_{\rm ad}\rangle_{\rm eff}=\langle\mathbbm{1}\rangle_{\rm eff}=1\neq\sum_{j}\chi^{*}_{j}\frac{\partial\theta^{*}_{j}}{\partial\theta_{i}}=0\;,$ (50)
since the identity can not influence the RG-flow as it has no influence on the Hamiltonian and therefore ${\nicefrac{{\partial\theta^{*}_{j}}}{{\partial\theta_{i}}}=0}$. In the bare Hilbert space, $\hat{\Omega}_{\rm ad}$ behaves the same as a regular observable, for which Eq. 43 holds. Thus, it follows from Eq. 46 that knowledge of ${\langle\hat{\Omega}_{\rm ad}\rangle_{\rm bare}}$ does not allow to establish a relationship between ${\langle\hat{h}_{i}\rangle_{\rm eff}}$ and ${\langle\hat{h}_{i}\rangle_{\rm bare}}$ through rescaling ${\langle\hat{h}_{i}\rangle_{\rm bare}}$ with ${\langle\hat{\Omega}_{\rm ad}\rangle_{\rm bare}}$. Hence, the importance of Eq. 30 comes not from accounting for scale differences, but from demonstrating that the optimization problem Eq. 8 is overdetermined when using the projectors $\hat{P}_{QS}$ as effective interactions and this holds at all scales due to Eq. 50.
V Results
Figure 3: Performance metrics for the optimized effective models. Here we compare the bare sAIM (points) with the results of F-learning (blue lines), and UML employing the minimal (green), 1e (orange) and 2e (red) models. As a figure of merit, panel a) shows the Kondo scale $T_{K}$ for these models, while b) gives the relative error of the effective $T^{\rm eff}_{K}$ with respect to the bare $T^{\rm bare}_{K}$. Panel c) shows the dependence of RG-relevant and -irrelevant operators on $U_{d}$. To illustrate the relative quality of the various effective models, in panel d) we show their impurity entropy (lines), compared with benchmark results for the sAIM (points), for representative bare $U_{d}=0.4$. Throughout we have used $0.38V^{2}/U_{d}=$ 0.3.
As proof of principle we only study mappings for the AIM, which can be solved completely using Wilson’s NRG (see Appendix C for the details of the calculations). We can use the analytical mapping to the minimal model Eq. 24 with the coupling $J_{SW}$ as comparison for our results. This mapping gives us also a reference value for the Kondo temperature
$T_{K}(D,J_{SW})=\alpha D\sqrt{\rho_{0}J_{SW}}\exp[-1/\rho_{0}J_{SW}+\gamma\rho_{0}J_{SW}]\;,$
with $\rho_{0}=1/2D$ the density of states of the bath at the Fermi level, $\gamma=\pi^{2}/4$ and $\alpha=\mathcal{O}(1)$. Our tests consist of performing the minimization of the KLD for the minimal model, the 1e-imp and the 2e-imp for an AIM with $8V^{2}/U_{d}=0.3D$ which leads to a constant $T_{K}(D,J_{SW})$. The interaction strengths are drawn from between $0.01\leq U_{d}/D\leq 100$ such as to cover the cases $U_{d}/D\ll 1$ and $U_{d}/D\gg 1$. The true Kondo temperature for the bare and effective models is extracted from the impurity entropy $S(T)$ [65], as shown in Fig. 3 d) for different optimized effective models and the corresponding bare model. In the following we us the definition $S(T_{K})=0.5$.
The first example is the minimization of the KLD for the minimal model Eq. 24, which contains only the RG-relevant spin-exchange term. The associated coupling $J$ gets strongly renormalized and as a result $\frac{\partial J^{*}}{\partial J}$ has a different functional form for bare and effective model depending on whether $U_{d}/D\ll 1$ or $U_{d}/D\gg 1$. In the case of $U_{d}/D\gg 1$ the AIM is effectively already in a LM fixed point and thus the flow of the AIM and the minimal model are coinciding, which means that because of matching the spin-spin exchange term the minimal model is also able to match $T_{K}$. This behaviour is clearly shown in Fig. 3 a) in green. When $U_{d}/D=\mathcal{O}(1)$ the free orbital fixed point plays a significant contribution in the flow of the AIM, which explains the behaviour in Fig. 3 a) as matching the observable is now no longer equal to matching $T_{K}$. Thus, the minimal model is not fit to yield good approximations of the $T^{\rm bare}_{K}$ based on knowledge of the observables. As a result the relative error $\Delta(T^{\rm eff}_{K},T^{\rm bare}_{K})$ diverges for $U_{d}\rightarrow 0$.
To understand more realistic scenarios only the parameter range $U_{d}/D<1$ is of importance, which is where the most improvement is needed over the minimal model. This can be achieved with the most general $\rm SU(2)$ symmetric Hamiltonian we can define on $\mathcal{H}^{\rm bath}_{0}\otimes\mathcal{H}^{\rm mol}_{\rm eff}$. However, keeping all projectors in Eq. 32 can not converge the KLD. At least one of the projectors must be eliminated. To help to choose which one we introduce the Frobenius scalar product (FSP) on the operator space $\langle\hat{A},\hat{B}\rangle=\textrm{tr}\left[\hat{A}^{\dagger}\cdot\hat{B}\right]$ and the corresponding norm $\|\hat{A}\|=\sqrt{\langle\hat{A}^{\dagger},\hat{A}\rangle}$. Using the FSP any projector $\hat{h}_{i}$ can be decomposed into its contribution from the FP Hamiltonian $\langle\hat{h}_{i},\hat{H}^{*}\rangle/\|\hat{H}^{*}\|^{2}=\alpha_{i}$ and an RG-irrelevant operator $\langle\hat{h}_{i},\hat{\Lambda}_{i}\rangle/\|\hat{\Lambda}_{i}\|^{2}=\beta_{i}$, with $\langle\hat{\Lambda}_{i},\hat{H}^{*}\rangle=0$. The operator $\hat{\Lambda}_{i}$ can easily be found after one round of the Gram-Schmidt procedure $\hat{\Lambda}_{i}=\hat{h}_{i}-\alpha_{i}\hat{H}^{*}$.
Figure 4: Panel a) and b) compare different configurations of Eq. 32 as effective Hamiltonians. The top left panel of a) shows the observables for the $\hat{H}^{JU}_{\rm eff}$ model, while the others show the observables of the $\hat{H}^{1\rm e-imp}_{\rm eff}$ for different eliminated interactions. The panels in a) are labelled by the eliminated interaction $\theta_{QS}=0$. Panel b) compares the relative error of $T_{K}$ for these four models.
The decomposition of the projector $\theta_{i}\hat{h}_{i}=\theta_{i}(\alpha_{i}\hat{H}^{*}+\beta_{i}\hat{\Lambda})=\gamma_{i}\hat{H}^{*}+\lambda_{i}\hat{\Lambda}_{i}$ can be used for a substitution
$\begin{gathered}\frac{\partial}{\partial\theta_{i}}\mathcal{F}(\{K\},\theta_{i})\Big{|}_{\theta_{i}=0}=\\ \Big{[}\beta_{i}\underbrace{\frac{\partial}{\partial\lambda_{i}}}_{\rm RG~{}irrelevant}+\alpha_{i}\underbrace{\frac{\partial}{\partial\gamma_{i}}}_{\rm RG~{}relevant}\Big{]}\mathcal{F}(\{K\},\lambda_{i},\gamma_{i})\Big{|}_{\gamma_{i}=\lambda_{i}=0}\;,\end{gathered}$ (53)
which tells us that the larger $\alpha_{i}$ the stronger $\langle\hat{h}_{i}\rangle$ flows under RG. For the example of the minimal model it is trivially $\beta_{i}=0$ and the only contribution to the flow is RG-relevant. The projectors $\hat{P}_{QS}$ (Eq. 32) on the other hand have a non-trivial decomposition. To render the overlap parameter $\alpha_{i}$ comparable we introduce $\tilde{\alpha}_{i}=|\langle\hat{h}_{i},\hat{H}^{*}\rangle/\|\hat{H}^{*}\|\|\hat{h}_{i}\||$, which is normalized with respect to the effective interaction. This provides a means to compare the strength of the renormalization of the effective interactions and therefore to decide which interaction to eliminate. For the effective interactions of $\hat{H}^{JU}_{\rm eff}$ and $\hat{H}^{1e-imp}_{\rm eff}$ the FSP with $\hat{H}^{*}=\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}$ is:
$\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}$ $\left(\hat{n}_{0}-1\right)^{2}$ $\hat{P}_{0,1}$ $\hat{P}_{0,0}$ $\hat{P}_{\pm 1,1/2}$
$\tilde{\alpha}\approx$ $1$ 00 $0.5$ $0.87$ 00
Based on the values of $\tilde{\alpha}_{i}$ the ideal Hamiltonian is $\hat{H}^{1e-imp}_{\rm eff}$ with $\theta_{0,0}=0$, as that excludes the term that gets renormalized the strongest. We put this hypothesis to the test by optimizing all four possible Hamiltonians, starting with $\hat{H}^{JU}_{\rm eff}$. In $\hat{H}^{JU}_{\rm eff}$ we have already set $L=0$, which means that the minimization of the KLD can be converged. As Fig. 4 a) shows, the observables can be matched for any $U_{d}$, however Fig. 4 b) shows that the relative error $\Delta(T^{\rm eff}_{K},T^{\rm bare}_{K})$ of the Kondo temperature is roughly equal to what can be obtained with the minimal model. This shows that the naive inclusion of RG-irrelevant terms does not improve the inference of energy scales, as is to be expected from Eq. 53. A very similar result in terms of $\Delta(T^{\rm eff}_{K},T^{\rm bare}_{K})$ is produced by optimizing $\hat{H}^{1e-imp}_{\rm eff}$ with $\theta_{1,0}=0$. However, Fig. 4 a) shows that the observables that get optimized match in bare and effective model for all $U_{d}$, while $\langle\hat{P}_{0,1}\rangle$ only matches if $U_{d}/D\rightarrow\infty$. That implies that in this limit $\langle\hat{\mathbb{I}}_{\rm eff}\rangle_{\rm eff}\rightarrow\langle\hat{\Omega}^{\dagger}_{\rm ad}\cdot\hat{\mathbb{I}}_{\rm eff}\cdot\hat{\Omega}_{\rm ad}\rangle_{\rm bare}$, which is clearly the case as double and zero-occupation states of the impurity are strongly suppressed. Also having $\theta_{\pm,1/2}=0$ does not improve $\Delta(T^{\rm eff}_{K},T^{\rm bare}_{K})$, but keeps it roughly at the same quality as the minimal model. Only when eliminating the strongest flowing observable $\theta_{0,0}=0$ the relative error seizes to diverge for $U_{d}/D\rightarrow 0$, as Fig. 4 a) shows. The results of minimizing the KLD with $\theta_{0,0}=0$ are shown as orange line in Fig. 3 a) which follows the bare $T_{K}$ curve orders of magnitude closer than the minimal model curve (green). The corresponding coupling parameters for the optimized 1e-imp model can be mapped to the parameters of Eq. 35. The coupling of the RG-relevant term can be isolated and compared to the coupling for the minimal model, which was tuned to reproduce the bare free energy [47]. Both couplings show a great correspondence, meaning that 1e-imp provides a good balance between capturing thermodynamic observables and energy scales. Just like for the minimal model the limit $U_{d}/D\gg 1$ yields the SW result where $J\rightarrow 0.3$ and thus, the RG irrelevant term $U_{0}\rightarrow 0$ vanishes.
Figure 5: Panel a) compares the 2e-imp effective Hamiltonian with (red) and without (purple) the off-diagonal terms Eq. 39. For the 2e-imp with the off-diagonal terms, panel b) shows the evolution of $T_{K}$ with respect to parameter update steps (epochs). The four panels of c) show the evolution of the relative error of observables with respect to the epoch. The observables are colour coded with the quantum numbers $(Q,S,T)$. Plotted for $U_{d}=0.4$.
This shows that the FSP provides a reliable way to identify ill-fitted interactions for the optimization. It also exemplifies that observables with at most a moderate flow ($\tilde{\alpha}\approx 0.5$) are best suited for the KLD distinguishability measure. Thus, 1e-imp with $\theta_{0,0}=0$ shows an enormous improvement over the observable-matched minimal model. We now show that the performance of the effective model can be systematically improved by including more RG-irrelevant terms in the form of the 2e-imp Hamiltonian.
The projectors of the 2e-imp effective Hamiltonian are constructed from the basis |$\ket{Q,S,T,s_{z};m}$ Appendix B, which are defined on $\mathcal{H}^{\rm bath}_{1}\otimes\mathcal{H}^{\rm bath}_{0}\otimes\mathcal{H}^{\rm mol}_{\rm eff}$. In this basis there are a total of 15 $s_{z}$-multiplets with several degenerate spin-charge subspaces and in particular a four-fold degeneracy of the subspace $Q=0,~{}S=1/2$, as shown in Eq. III. Transitions between multiplets are not prohibited by symmetries. Thus not including such transitions means limiting the expressiveness of the model, while however guaranteeing a convex loss function. The impact of including the off-diagonal terms Eq. 39 can be quantified using the FSP. An optimized representation of any operator $\hat{X}$ using $\hat{H}^{n\rm-mol}_{\rm eff}(\{\theta\})$ can be found by minimizing the normalized FSP [66]
($\mathcal{L}(\{\theta\},\hat{X})=1-\frac{\left|\Big{\langle}\sum_{i}\theta_{i}\hat{P}_{i},\hat{X}\Big{\rangle}\right|}{\Big{\|}\sum_{i}\theta_{i}\hat{P}_{i}\Big{\|}\cdot\|\hat{B}\|}\;.$ (54)
In the following optimization of Eq. 54 are carried out using vanilla gradient descent. To satisfy $T^{\rm bare}_{K}\approx T^{\rm eff}_{K}$ in the limit $U_{d}/D\gg 1$ the effective Hamiltonian must be able to represent the Kondo Hamiltonian Eq. 24. With $\hat{X}=\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}$ it can be determined that including the off-diagonal terms optimizing Eq. 54 yields ($\mathcal{L}(\{\theta_{\rm op}\},\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0})=0$, whereas without the off-diagonal terms the representation is limited to ($\mathcal{L}(\{\theta_{\rm op}\},\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0})\approx 0.3$. This is an enormous difference in expressiveness and since the $U_{d}/D\gg 1$ limit is important to satisfy, we prioritise generality over convexity in the case of 2e-imp and include the off-diagonal elements. Note that choosing $\hat{T}$ as a defining operator for the 2e-imp basis means that with and without off-diagonal elements we can find $\{\theta_{\rm op}\}$ such that ($\mathcal{L}(\{\theta_{\rm op}\},\hat{T})=0$. Including the off-diagonal terms leads to a total of 25 interactions in the 2e-imp Hamiltonian. Again, the interaction with the largest $\tilde{\alpha}$ is omitted from the Hamiltonian. The minimization of Eq. 8 can still be carried out using GD, however local minima can be encountered. We have found that running in to local minima can be avoided by initializing the optimization with the expectation values of the bare model $\langle\hat{P}_{QS;t,t^{\prime}}\rangle_{\rm bare}\xrightarrow{\rm initialize}\theta_{QS;t,t^{\prime}}$. This has the advantage that symmetries of the model are automatically respected and small bare expectation values lead directly to small effective parameters.
We have established that 2e-imp model is with the off-diagonal terms capable of representing every possible real $\rm SU(2)$ symmetric Hamiltonian on the effective Fock space. However, it is incapable to represent all possible $\rm SU(2)$ symmetric Hamiltonians on the bare Fock space. As a result we find that minimizing the KLD is only approximately possible, since there are observables that can not be simultaneously matched, despite that one observables has been omitted. We give an example for this in Fig. 5 c), where the relative error of the expectation values of all diagonal projectors is shown for every step of the optimization. The figure shows that most observables converge exponentially fast, while some are seemingly stationary. While this means that the optimization can not reach the minimum of the KLD, the parameters however converge and so does the $T_{K}$ as seen in Fig. 5 b). The results of the minimizations for all $U_{d}$ are shown in Fig. 3, we find an improvement of the relative error of the Kondo temperature (Fig. 3 b) ) of an order of magnitude with respect to the 1e-imp result in the $U_{d}/D\ll 1$ regime. This result can be compared to the convex 2e-imp optimization problem without off-diagonal terms as done in Fig. 5 b). As expected the more general effective Hamiltonian performs better for $U_{d}/D\gg 1$, however the less general Hamiltonian yields improved accuracy at $U_{d}/D\ll 1$. This can be attributed to the off-diagonal elements that include terms that are more RG-relevant, which is clearly necessary to perform well at $U_{d}/D\gg 1$ but in the $U_{d}/D\ll 1$ regime these observables are stronger renormalized and hence impact the performance. However, the impact on the performance of 2e-imp is small compared to the improvement over 1e-imp, as shown in Fig. 5 b).
An other metric that indicates how well an effective model captures the bare model are observables that get not explicitly matched during the optimization process. The spin exchange term is very interesting in that regard, as it is the only RG-relevant term in the system and thus experiences the strongest renormalization. As expected, Fig. 3 c) shows that at $U_{d}/D\gg 1$ all effective models coincide with the bare model. For $U_{d}/D\ll 1$ the minimal model is trivially still perfectly matching the bare model. However, we find that while 1e-imp performs worse than the $\mathcal{Z}$-matched minimal model, the 2e-imp greatly improves over both. An other insightful observable is the density operator on the 00-orbital $(1-\hat{n}_{0})^{2}$, which is RG-irrelevant and does only get weakly renormalized. In fact 2e-imp and the $\mathcal{Z}$-matched minimal model reproduce the bare value in the $U_{d}/D\gg 1$ and $U_{d}/D\ll 1$ regimes. The most significant deviations occur at $U_{d}/D\approx 1$, they are however still small. The advantage of this analysis is that other than the $T_{K}$, observables are necessarily computable without solving the entire bare model and thus, allow cross comparison with $\Delta(\langle\hat{h}_{i}\rangle_{\rm eff},\langle\hat{h}_{i}\rangle_{\rm bare})$.
Figure 6: Panel a) shows the evolution of optimized parameters of $\hat{H}^{1\rm e-imp}_{\rm eff}$ with temperature for the sAIM. The corresponding observables for the sAIM and optimized effective model are shown in b). The temperature axis is logarithmic, but includes a cut to $T=0$ to indicate that optimization can be done using only ground state properties if required.
This shows that energy scales and general thermodynamic observables can be captured by matching only a specific set of observables. The quality of the approximation can be improved systematically by incorporating more RG-irrelevant terms in the effective Hamiltonian.
[…]

### v2

@@ IV.3 UML optimization of extended models
When learning a representation for complex bare models (for example ab initio models of a realistic molecular junction), the Kondo temperature $T_{K}$ will not typically be a good figure of merit to assess the quality of the effective model. This is because, for any practical application of UML, the bare model will not itself be solvable down to very low but finite temperatures – hence the need in the first place for a simplified effective model that can be fully solved. Therefore one will not in general have access to the reference $T_{K}$ for the bare model, from which to compute the relative error Eq. 51. The error in computable observables is thus more practical; the KLD loss function gradient depends on these observables, which are matched during the optimization process. For simple bare model test cases such as the AIM, we can however still compare with $T_{K}$ and other quantities.
Taking the AIM again as our bare model for demonstration, we perform UML optimization on the extended MC effective model Eq. 61. For representative AIM parameters, in Fig. 6(a) we plot the relative error of selected (diagonal) observables labelled by the projector quantum numbers $(Q,S,T)$, as a function of the optimization step (epoch). We find that most effective model observables converge exponentially quickly to their reference bare model counterparts, whereas a few are essentially stationary and cannot be matched. However, as Fig. 6(b) shows, this does not make much difference to the quality of the optimized effective model, as quantified by the Kondo temperature $T_{K}$. After $\sim 100$ or so training steps, $T_{K}$ for the effective model (red line) converges quite accurately to that of the bare model (blue line).
In Fig. 6(c) we show the relative (logarithmic) error in $T_{K}$ in the fully effective model, as a function of AIM interaction strength $U_{d}$, with the red line for the fully general extended MC effective model Eq. 61, compared with the fixed-basis model Eq. 59 as the purple line. We see that $T_{K}$ is reproduced rather well by the effective model in both cases, for all values of $U_{d}$. As expected from the above arguments, the generalized model does better at large $U_{d}$. However, somewhat unexpectedly, the fixed-basis model has a smaller error at small $U_{d}$. We attribute this to the fact that optimization of the fixed-basis model requires to match fewer observables, which is easier to achieve in practice. Furthermore, the additional (off-diagonal) observables in Eq. 61 appear to flow more strongly under RG at small $U_{d}$ than the projectors in Eq. 59, meaning that matching them may ironically reduce the quality of the optimized effective model. Despite these subtleties of the training process, the extended MC effective models do represent a systematic improvement over simpler effective models.
We compare the different approaches in the following.
V Comparison of models
Figure 7: Comparison of performance metrics for the various optimized effective models considered in this work, with the AIM as the benchmark bare model. Results obtained by NRG. (a) Kondo temperature $T_{K}$ as a function of $U_{d}$, comparing the AIM (points) and $\mathcal{F}$-learned Kondo result (blue line) with the UML-optimized effective models: minimal (green), MC (orange), and extended MC (red). Corresponding $T_{K}$ relative error shown in (b) and normalized observables shown in (c). (d) Impurity contribution to the entropy $S_{\rm imp}(T)$ vs $T/U_{d}$ for the bare AIM with $U_{d}=0.4$ (points) and the optimized effective models (lines). In all cases we set $0.38V^{2}/U_{d}=$ 0.3.
In this section we consider the relative performance of the various effective models introduced in this paper, optimized by UML to best describe the the physics of the AIM as our benchmark bare model. As we have shown, it is in general not possible to simultaneously capture within a simplified effective model the universal low-temperature thermodynamics (as characterized by the Kondo temperature $T_{K}$) as well as the behavior of static thermal expectation values of local operators. Any effective model must therefore strike a compromise in reproducing these features. Different methods may target different aspects of the physics and be better suited to different applications.
The $\mathcal{F}$-learning approach [48] requires calculation of the full free energy of bare and effective models within the optimization process. It guarantees to reproduce the low-energy thermodynamics and $T_{K}$ within a minimal Kondo model, but fails to capture local observables except in the perturbative limit $U_{d}\gg D$ where the SWT also holds as a good description. However, it might not be possible to accurately calculate the free energy of a complex bare model in practice, so other methods might be preferred.
UML, by contrast, is based on matching certain local observables, and so by construction these observables agree in bare and optimized effective models. However, applied to crude effective models, such as minimal Kondo-type models that only feature the most RG-relevant terms, UML yields solutions that do not capture well the low-energy thermodynamics and Kondo scale $T_{K}$. By including minimally constrained models in an expanded effective impurity Fock space, we systematically improve upon results obtained by UML. Such models include RG-marginal and RG-irrelevant corrections, and UML optimization is done using the set of observables that do not flow strongly. Depending on the level of theory employed, one can optimize with UML generalized effective models that do capture local observables while also approximating $T_{K}$ with good accuracy.
These features are illustrated in Fig. 7, where we compare the performance of the different effective models.
The UML method is based on minimizing the KLD loss function Eq. 9 that quantifies the distinguishability between bare and effective model. In general ML applications, the value of the loss function is a useful metric to judge the training progress and ultimately the performance of the model. In the case of Eq. 9 however, the KLD is essentially impossible to compute in practice. The training progress cannot therefore be established in the standard way. On the other hand, we have shown that to minimize the KLD via GD we only need to know its gradient, Eq. 10. This can be done without actually ever evaluating the KLD itself. For a commuting set of local impurity operators (for example the symmetry projectors in MC models), the KLD is convex and so the gradient vanishes only at saddle-points or global minima of the KLD. Thus, the gradient is itself a good replacement for the loss function to determine whether the training of a model has converged. For the cases considered in Fig. 7, the KLD gradient can always be made to vanish.
We therefore turn to our other metrics to quantify the relative performance of the models. With our NRG solution of the bare and optimized effective models, we can compute the Kondo temperature directly. This is shown in Fig. 7(a) as a function of $U_{d}$. The exact result for the bare AIM (points) are compared with the effective models (lines): the $\mathcal{F}$-learning and UML result for the minimal Kondo model shown as blue and green lines; the MC and extended MC models shown as the orange and red lines. Panel (b) shows the corresponding relative error $\Delta(T^{\rm eff}_{K},T^{\rm bare}_{K})$ evaluated using Eq. 51. In panel (c) we plot the $T=0$ value of static observables $\langle\hat{O}\rangle=\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle$ and $\langle(\hat{n}_{0}-1)^{2}\rangle$, comparing results in bare AIM (points) with those of the optimized effective models (lines). We have normalized the results by their value in the $U_{d}/D\to\infty$ limit.
First, we note that by construction, $\mathcal{F}$-learning perfectly reproduces the $T_{K}$ value for all $U_{d}$, blue line in Fig. 7(a). However the behavior of $\langle\hat{O}\rangle=\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle$, which is not directly optimized by $\mathcal{F}$-learning, is poorly captured, see panel (c). On the other hand UML for the minimal Kondo model (green line) directly matches the observable $\langle\hat{O}\rangle=\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle$ in panel (c) and therefore by construction agrees perfectly with the bare AIM results. The $T_{K}$ in this case is not directly matched, and panel (a) shows that UML for the minimal model performs very poorly using that metric, except in the SWT (perturbative) limit $U_{d}/D\to\infty$.
The behavior of the Kondo temperature and the observables is a stringent nontrivial performance check for the MC and extended MC models, since neither $T_{K}$ nor $\langle\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}\rangle$ are directly matched within UML. We find that UML represents a good compromise on the performance of these models for the two metrics. Good performance is seen for each over all $U_{d}$, with a systematic improvement for extended MC over plain MC.
With the optimized effective models at hand, we compute using NRG the full temperature-dependence of the thermodynamic entropy for a representative example of the bare AIM with $U_{d}=0.4D$, comparing the results in Fig. 7(d). Within the regime of validity of the effective models ($T/U_{d}\ll 1$), we see that the MC and extended MC models do accurately capture the physics of interest for the AIM.
VI Conclusion
[…]